# CAGE: Curvature-Aware Gradient Estimation For Accurate Quantization-Aware Training

### v1

@@ Abstract
Despite significant work on low-bit quantization-aware training (QAT), there is still an accuracy gap between such techniques and native training. To address this, we introduce CAGE (Curvature-Aware Gradient Estimation), a new QAT method that augments the straight-through estimator (STE) gradient with a curvature-aware correction designed to counteract the loss increase induced by quantization. CAGE is derived from a multi-objective view of QAT that balances loss minimization with adherence to quantization constraints, yielding a principled correction term that depends on local curvature information. On the theoretical side, we introduce the notion of Pareto-optimal solutions for quantized optimization, and establish that CAGE yields strong convergence guarantees in the smooth non-convex setting. In terms of implementation, our approach is optimizer-agnostic, but we provide a highly-efficient implementation that leverages Adam statistics. When pre-training Llama-style models of up to 800M-parameters, CAGE recovers over 10% of the quantization-induced loss increase in the W4A4 regime over outlier-mitigation methods. These results indicate that curvature-aware gradient corrections can bridge the remaining performance gap beyond current outlier-handling methods.
@@ 1 Introduction
Quantization has emerged as a standard technique for improving the computational efficiency of large language model (LLM) deployments, as prominent open-source models such as Llama, Gemma, and GPT-OSS are released in open-source formats [30; 29; 25]. Yet, the vast majority of research in this area has concentrated on post-training quantization (PTQ), where a fully trained model’s weights are quantized by applying numerical algorithms over a small calibration dataset [13].
@@
Relatively less is known about quantization-aware training (QAT), where the model learns to adapt to the constraints of quantization during the training process itself. QAT is more computationally-expensive, but can yield better accuracy. The state of the art for QAT methods has long been dominated by the straight-through estimator (STE), a method first suggested by Hinton [15] and formalized by Bengio et al. [5], integrated by default in deep learning frameworks like PyTorch [28]. This approach bypasses the non-differentiable quantization function in the backward pass, by approximating its gradient with the identity. While versatile, STE is known to suffer from slow convergence and instability. A body of work, such as LSQ [12], LSQ+ [6], and the recently proposed QuEST [11; 6; 26], has introduced improved approaches to stabilize and accelerate this process learnable quantization parameters or more sophisticated gradient scaling. Yet, these methods are heuristic, and do not offer formal convergence guarantees.
Contribution.
@@
In this paper, we address this gap by examining QAT from both a theoretical and a practical standpoint. We reframe QAT as a multi-objective optimization problem, where the goal is to simultaneously minimize the task loss and the quantization error. From this perspective, we identify a condition for Pareto-optimality, where any improvement in one component of the objective necessarily leads to a degradation in the other. This formulation inspires a new algorithm, called CAGE, for Curvature-Aware Gradient Estimation, which provides theoretical guarantees with high accuracy and efficiency in practice. More precisely, our technical contributions are as follows:
@@
  • The CAGE we introduce provides a principled method for augmenting the standard STE gradient with a curvature-aware correction term, derived directly from our Pareto-optimality condition. This explicitly counteracts the increase in loss induced by the quantization step, by leveraging local second-order information about the loss landscape (i.e., the Hessian).
@@
  • From the theoretical perspective, we prove that CAGE possesses strong ergodic convergence guarantees to a Pareto-optimal point in the smooth non-convex setting, directly addressing the lack of guarantees that is a primary limitation of prior work. Our approach is optimizer-agnostic, and we provide a highly efficient implementation that utilizes statistics readily available in adaptive optimizers such as Adam.
  • We validate the effectiveness of CAGE through extensive experiments, including pre-training Llama-style models of up to 800M parameters from scratch. Our results demonstrate that in the challenging low-bit W4A4 (4-bit weights and 4-bit activations) regime, CAGE recovers over 10% of the quantization-induced loss increase when compared to strong baselines that incorporate modern outlier-mitigation techniques.
@@
In summary, our findings show that incorporating curvature-aware gradient corrections is a key step toward bridging the remaining performance gap between low-bit quantized models and their full-precision counterparts.
2 Related Work
@@
The central challenge in QAT is the non-differentiable quantization function, with zero gradients almost everywhere. The standard is the STE, initially suggested by Hinton [15] and formalized by Bengio et al. [5], which bypasses the quantization operator during the backward pass. Yet, STE is known to often lead to instability and suboptimal convergence [32].
@@
Thus, substantial research has focused on STE refinements, via e.g. learnable quantization parameters, e.g. [12], or improving the accuracy of the gradient itself, such as AdaSTE [20] (working via complex bi-level optimization), or ReSTE [31] which uses a rectified estimator which balances the estimation error and the gradient stability. In addition, ProxQuant [3] reformulated QAT as a regularized learning problem solved via proximal gradient methods. Instead of relying solely on STE, ProxQuant applies a prox operator between stochastic gradient steps to encourage quantization. Liu et al. [23] proposed Reinmax, based on the observation that the STE acts as a first-order approximation of the gradient and proposed a second-order estimator via Heun’s method. In addition, Markov Random Field (MRF) and lifted probability space formulations [1] represent each parameter of the model as weighted average of quantization nodes and minimizes the loss with respect to these probabilistic weights.
The Mirror Descent formulation interprets full-precision parameters as the dual of quantized ones and analyses the cases where quantization operator generates a valid mirror map. Particularly, if we take softmax projection as quantization operator, then the dual iterates of STE follow Exponentiated Gradient Descent (EGD) [2]. By contrast, CAGE introduces a new approach rooted in a multi-objective optimization perspective of QAT. Instead of refining the first-order approximation or employing standard regularization, CAGE introduces a principled, curvature-aware correction term to the gradient, with theoretical guarantees.
@@ 3.1 Problem Description and Motivation
Intuitively, the key question in quantization-aware training (QAT) is to construct a gradient estimator that updates the parameters towards the an “optimal” quantized model, i.e. a constrained model which minimizes the loss. More precisely, the optimization problem underlying QAT can be described by the following constrained optimization:
@@
where $\mathcal{Q}=\{x\in\mathbb{R}^{d}\colon x=Q(x)\}$ is the constraint set of quantized parameters and $Q\colon\mathbb{R}^{d}\to\mathbb{R}^{d}$ is the quantization operator. An alternative way of looking at the same problem is to consider its unconstrained reformulation and minimizing the objective $f(Q(x))$ with respect to $x\in\mathbb{R}^{d}$ [21]. The key issue here is the non-differentiability of quantization operator $Q$. As such, the straight-through estimator [4] is the default choice to approximate the gradient of $f(Q(x)$. Specifically, $\nabla_{x}[f(Q(x))]=JQ(x)^{\top}\cdot\nabla f(Q(x))$ is approximated by $\nabla f(Q(x))$. Effectively, this replaces the Jacobian of the quantization operation by the identity matrix [7; 17; 21].
@@
Error feedback reformulation. One way to gain more insight into the dynamics of STE is to examine it through the lens of error feedback. It is easy to observe that SGD-based training with STE over the iterate $x_{t}$ is equivalent to an instance of error-feedback [8] with quantized parameters $w_{t}=Q(x_{t})$ and quantization error $e_{t}=x_{t}-Q(x_{t})$, as described over iterations $t\geq 0$ below:
$x_{t+1}=x_{t}-\alpha\widetilde{\nabla}f(Q(x_{t}))\quad\iff\quad\begin{cases}g_{t}=\alpha\widetilde{\nabla}f(w_{t})-e_{t}\\ w_{t+1}=Q(w_{t}-g_{t})\\ e_{t+1}=(w_{t}-g_{t})-w_{t+1}.\end{cases}$
@@
Effectively, the equivalent right-hand-side formulation says that STE is equivalent to a process where the quantization error $e_{t+1}$ due to weight quantization at a given step $(w_{t}-g_{t})-w_{t+1}$ is fed into the gradients at the next step. This would be “correct” if the loss function $f$ were an isotropic quadratic function, but is not the right update in general.
@@
Multi-objective perspective. Prior work on convergence guarantees for the problem (1) in the non-convex setting attempt to find quantized point $Q(x^{*})$ that is a stationary point for the loss $f$, namely $\nabla f(Q(x^{*}))=0$. However, as quantization is not an invertible transformation, this goal cannot be achieved in general, and current convergence guarantees have non-vanishing terms in their rate, proportional to quantization error [22; 8; 27]. For the sake of illustration, consider a toy example where our scalar loss function is $f(x)=\frac{1}{2}(x-\frac{1}{2})^{2}$, and the quantization operator $Q(x)=\lfloor x\rfloor$ gives the integer part of the one dimensional input $x\in\mathbb{R}$. Clearly, $\nabla f(Q(x))=\lfloor x\rfloor-\frac{1}{2}$ does not vanish at any point, with the smallest absolute value being $\frac{1}{2}$, coming from the quantization/rounding error.
@@
Instead, we propose to view the QAT problem in Equation (1) from the perspective of multi-objective optimization. The first objective is to minimize the loss $f(x)$ (in non-convex setting this becomes finding a stationary point, namely solving $\nabla f(x)=0$), while the second is to satisfy quantization constraints $x\in{\mathcal{Q}}$. In other words, we aim to find a solution $x^{*}$ for which $\nabla f(x^{*})=0$ and the distance between $x^{*}$ and $Q(x^{*})$ is minimized. As noted before, it is easy to see that, these two conditions do not have to hold in general, meaning $\nabla f(x)$ can be non-zero for any quantized point $x\in{\mathcal{Q}}$. Instead, we define solutions as Pareto optimal points that balance these two conditions.
@@
Let $x\in\mathbb{R}^{d}$ be the current state of parameters. If we update towards $\nabla f(x)$ (or any other direction that aligns with it) with small enough step-size, then we would minimize the loss. Clearly, the opposite direction would maximize the loss. Analogously, if we update towards Q(x)xQ(x)-x, then we would reduce quantization error. To have those two objectives in balance, we want our iterates to converge to some Pareto-optimal state $x^{*}$ such that $\nabla f(x^{*})=\lambda(Q(x^{*})-x^{*})$ for some scalar $\lambda>0$. In this case, any sufficiently small update applied to $x^{*}$ would hurt at least one of the objectives. In other words, any local improvement of one objective would be at the cost of other objective. Therefore, we define $\lambda$-Pareto optimal solution $x^{*}$ for the problem (1) as
@@
where $\lambda>0$ is a parameter balancing the two objectives. Recalling the toy example mentioned earlier, the value $x^{*}(\lambda)=\frac{1}{2(1+\lambda)}$ is $\lambda$-Pareto optimal for any $\lambda>0$. Note that there can be many Pareto optimal solutions for the same $\lambda$. For example, $x^{*}=-\nicefrac{{1}}{{4}}$ is another $1$-Pareto optimal value.
Motivated by this multi-objective viewpoint and the new optimality condition in Equation (2), we propose incorporating the quantization error $e_{t}=x_{t}-Q(x_{t})$ directly into the training dynamics. By doing so, we implicitly regularize the initial task loss, which can be made explicit if the quantization scheme is smooth (as we demonstrate in the theoretical analysis). Indeed, under the smoothness assumption of $Q$ (Assumption 3), the quantization error can be expressed as $x-Q(x)=\nabla\phi(x)$ for some smooth regularizer $\phi(x)$. Notably, our approach naturally avoids the need to handle the (non-existent) Jacobian $JQ(x)$ of the quantization operator and bypasses the computation of a proximal step for the explicit “quadratic” regularizer $\frac{\lambda}{2}\|x-Q(x)\|_{2}^{2}$, which would be infeasible for the dynamic quantization used in our experiments [3].
@@
If the base optimizer is SGD, then these two corrections are identical and we provide theoretical convergence guarantees to a Pareto-optimal point in the smooth non-convex setting. However, for statefull optimizers, such as Adam, these two corrections produce conceptually different schemes, which we discuss below (subsuming bias correction terms into the learning rate $\alpha$):
$\textnormal{ Mini-batch gradient: }\quad\;g_{t}$ $=$ $\widetilde{\nabla}f(x_{t}){\color[rgb]{0,0,0}\;+\;\lambda_{t}e_{t}},\;{\color[rgb]{0,0,0}\textrm{(coupled correction)}}$
$\textnormal{ Optimizer states: }\quad m_{t}$ $=$ $\beta_{1}m_{t-1}+(1-\beta_{1})g_{t},\;v_{t}=\beta_{2}v_{t-1}+(1-\beta_{2})g_{t}^{2},$
$\textnormal{ Optimizer update: }\quad\Delta_{t}$ $=$ $m_{t}/(\sqrt{v_{t}}+\varepsilon){\color[rgb]{0,0,1}\;+\;\lambda_{t}e_{t}},\;{\color[rgb]{0,0,1}\textrm{(decoupled correction)}}$
$\textnormal{ Model update: }\;x_{t+1}$ $=$ $x_{t}-\alpha\Delta_{t},$
To compare these two correction variants in the context of the Adam optimizer, note that the coupled correction term $\color[rgb]{0,0,0}\lambda_{t}e_{t}$} becomes $(1-\beta_{1}){\color[rgb]{0,0,0}\lambda_{t}e_{t}}/(\sqrt{v_{t}}+\varepsilon)$ after being processed by the optimizer states. In contrast to the decoupled correction, the coupled version introduces diagonally-preconditioned curvature-aware correction term, where the curvature is approximated using the Adam statistics already maintained within the optimizer states.
@@
Concurrent work [18] proposes an approach called LOTION, which smooths the quantized loss using unbiased randomized rounding (RR) via $f(Q(x))\approx\mathbb{E}_{\epsilon\sim{\rm RR}}[f(x+\epsilon)]$, and considers its second-order expansion as a surrogate loss to minimize, i.e., $f(x+\epsilon)\approx f(x)+\epsilon^{\top}\nabla f(x)+\tfrac{1}{2}\epsilon^{\top}{\bf H}(x)\epsilon$. Due to the unbiasedness of rounding (i.e., $\mathbb{E}[\epsilon]=0$), the first-order term of the expansion vanishes, and the smoothed loss becomes the original loss $f(x)$ with an additional curvature-aware regularization term $\tfrac{1}{2}{\rm Tr}\left({\bf H}(x)\,{\rm Cov}[\epsilon]\right)$, where the exact Hessian matrix is replaced by its Gauss–Newton approximation. If we further apply the empirical Fisher approximation and approximate the diagonal entries of the Gauss–Newton matrix using Adam statistics, this regularization corresponds to our coupled correction variant. However, computing the gradient estimator for the regularized loss requires differentiating the regularizer, which involves third-order derivatives of the loss.
In contrast to LOTION, we propose to regularize the training dynamics directly rather than the loss itself. The theoretical advantage of our approach is that the regularized dynamics naturally arise from our Pareto-optimality condition, while the practical benefit is that it avoids the need to differentiate or compute the proximal operator of a loss regularizer. In addition, LOTION requires a full-precision forward pass, and the final model is a weight-only quantized model, while CAGE training is done with weights and activations quantized, where it can also benefit from high-performance low-precision GEMM kernels for a more efficient training and inference.

### v2

@@ Abstract
Despite significant work on low-bit quantization-aware training (QAT), there is still an accuracy gap between such techniques and native training. To address this, we introduce CAGE (Curvature-Aware Gradient Estimation), a new QAT method that augments the straight-through estimator (STE) gradient with a curvature-aware correction designed to counteract the loss increase induced by quantization. CAGE is derived from a multi-objective view of QAT that balances loss minimization with the quantization constraints, yielding a principled correction term that depends on local curvature information. On the theoretical side, we introduce the notion of Pareto-optimal solutions for quantized optimization, and establish that CAGE yields strong convergence guarantees in the smooth non-convex setting. In terms of implementation, our approach is optimizer-agnostic, but we provide a highly-efficient implementation that leverages Adam statistics. CAGE significantly improves upon the prior state-of-the-art methods in terms of accuracy, for similar computational cost: for QAT fine-tuning, it halves the compression accuracy loss relative to the prior best method, while for QAT pre-training of Llama models, its accuracy for 3-bit weights-and-activations (W3A3) matches the accuracy achieved at 4-bits (W4A4) with the prior best method. The official implementation can be found over https://github.com/IST-DASLab/CAGE.
footnotetext: *Equal contribution 1ISTA 2Red Hat AI. Correspondence to: Dan Alistarh <dan.alistarh@ist.ac.at>, Soroush Tabesh <soroush.tabesh@ist.ac.at>.
Copyright 2025 by the author(s).
@@ 1 Introduction
Quantization has emerged as a standard technique for improving the computational efficiency of large language model (LLM) deployments, as prominent open-source models such as Llama, Gemma, and GPT-OSS are released in compressed formats touvron2023llama2openfoundation; team2024gemma; openai:gptoss. Yet, the vast majority of open research in this area has concentrated on post-training quantization (PTQ), where a fully trained model’s weights are quantized by applying numerical algorithms over a small calibration dataset frantar2022gptq.
@@
Relatively less is known about quantization-aware training (QAT) bengio2013estimating; krishnamoorthi2018quantizing; esser2020lsq, where the model learns to adapt to the constraints of quantization during the optimization process itself. QAT is more computationally-expensive, but can yield better accuracy. The state of the art for QAT methods has long been dominated by the straight-through estimator (STE), a method first suggested by Hinton2012 and formalized by bengio2013ste, which is integrated by default in deep learning frameworks like PyTorch pytorch. This approach bypasses the non-differentiable quantization function in the backward pass, by approximating its gradient with the identity. While versatile, STE is known to suffer from slow convergence and instability. A body of work, such as LSQ (esser2020lsq), LSQ+ (bhalgat2020lsqplus), and more recent methods esser2019learned; bhalgat2020lsqplus; nagel2022overcoming; lee2021network, has introduced improved approaches to stabilize and accelerate this process learnable quantization parameters or more sophisticated gradient scaling. Yet, these methods are heuristic, and do not offer formal convergence guarantees.
Contributions.
@@
In this paper, we address this gap by examining QAT from both a theoretical and a practical standpoint. We provide a formal framework for QAT as a multi-objective optimization problem, where the goal is to simultaneously minimize the task loss and the quantization error. From this perspective, we identify a condition for Pareto-optimality, where any improvement in one component of the objective necessarily leads to a degradation in the other. This formulation inspires a family of algorithms, called CAGE, for Curvature-Aware Gradient Estimation, which provides both theoretical guarantees, and high accuracy and efficiency in practice. More precisely, our contributions are as follows:
@@
  • The CAGE method we introduce provides a principled method for augmenting the standard STE gradient with a curvature-aware correction term, derived directly from our Pareto-optimality condition. This term explicitly counteracts the increase in loss induced by the quantization step, by leveraging local second-order information about the loss landscape (i.e., the Hessian).
@@
  • From the theoretical perspective, we prove, under assumptions, that CAGE possesses strong ergodic convergence guarantees to a Pareto-optimal point in the smooth non-convex setting, directly addressing the lack of guarantees that is a primary limitation of prior work in this area. Our approach is optimizer-agnostic; we provide a highly-efficient implementation of this framework that utilizes statistics readily available in adaptive optimizers such as Adam.
  • We validate the effectiveness of CAGE through extensive experiments, including synthetic experiments meant to validate the theory, post-training fine-tuning of standard models, and pre-training of Llama-style models of up to 800M parameters from scratch. Our results demonstrate that CAGE leads to state-of-the-art results across all scenarios: for fine-tuning QAT, it roughly halves the quantization error relative to the prior best method, whereas, in pre-training experiments, CAGE-trained models with 3-bits weights and activations (W3A3) offer lower loss than 4-bit (W4A4) models trained with the prior best method (QuEST (panferov2025quest)).
  • Moreover, we show experimentally that the CAGE approach is optimizer-agnostic, as it leads to consistent gains across several optimizers, including AdamW (loshchilov2018adamw), Muon (muon) and Shampoo shampoo.
@@
In summary, we show that incorporating curvature-aware gradient correction via CAGE is a promising step toward bridging the remaining performance gap between low-bit quantized models and their full-precision counterparts.
2 Prior Work
@@
The central challenge in QAT is the non-differentiable quantization function, with zero gradients almost everywhere. The standard is the STE, initially suggested by Hinton2012 and formalized by bengio2013ste, which bypasses the quantization operator during the backward pass. Yet, STE is known to often lead to instability and suboptimal convergence (Yin2019).
@@
Thus, substantial research has focused on STE refinements through learnable quantization parameters (esser2020lsq), improving the accuracy of the gradient itself (Le2021AdaSTE) or incorporating loss regularization with proximal mapping (bai2018proxquant). For instance, AdaSTE (Le2021AdaSTE) enhances gradient estimation by viewing QAT problem from the perspective of bi-level optimization. ReSTE (Wu2023ReSTE) balances the estimation error and the gradient stability via a rectified estimator. liu2023bridging proposed Reinmax, noticing that the STE acts as a first-order approximation of the gradient and proposed a second-order estimator via Heun’s method.
In addition, ProxQuant (bai2018proxquant) reformulates QAT as a regularized learning problem solved via proximal gradient methods. Instead of relying solely on STE, ProxQuant applies a prox operator between stochastic gradient steps to encourage quantization. Similar to ProxQuant, PARQ uses convex and piecewise-affine regularization to induce weight quantization with the difference that gradient updates are added to the full-precision parameters before applying the proximal operator jin2025PARQ.
The Mirror Descent formulation interprets full-precision parameters as the dual of quantized ones and analyses the cases where quantization operator generates a valid mirror map. Particularly, if we take softmax projection as quantization operator, then the dual iterates of STE follow Exponentiated Gradient Descent (EGD) Ajanthan2021Ajanthan. Markov Random Field (MRF) and lifted probability space formulations Ajanthan2019MRF represent each parameter of the model as weighted average of quantized values, and minimizes the loss with respect to these probabilistic weights.
lee2021elementwisescaling proposed the element-wise gradient scaling (EWGS) approach, in which the scaling of each gradient element depends on both its sign and the corresponding quantization error. Additionally, these scalings are adjusted by a global scaling factor obtained from approximate Hessian information. A related method employs position-based scaled gradients, where the gradient estimator is adjusted only by multiplying the quantization error with a small positive margin kim2020positionbasedscaling.
Another phenomenon observed in STE-based training is the oscillation of weights around decision boundaries between adjacent quantization grid points défossez2022diffmodelcompression. To mitigate these oscillations and stabilize training, nagel2022oscillationsQAT introduced a regularization term on top of the loss to dampen the oscillations or alternatively proposed mechanisms for detecting oscillatory weights and freezing them during training.
By contrast to prior work, CAGE provides a multi-objective optimization perspective for QAT. Instead of refining the first-order approximation or applying standard regularization, CAGE introduces a curvature-aware correction term to the gradient.
@@ 3.1 Problem Description and Motivation
Intuitively, the key question in quantization-aware training (QAT) is to construct a gradient estimator that updates the parameters towards an “optimal” quantized model, i.e., a constrained model which minimizes the loss. More precisely, the optimization problem underlying QAT can be described by the following constrained optimization:
@@
where $\mathcal{Q}=\{x\in\mathbb{R}^{d}\colon x=Q(x)\}$ is the constraint set of quantized parameters and $Q\colon\mathbb{R}^{d}\to\mathbb{R}^{d}$ is the quantization operator. An alternative way of looking at the same problem is to consider its unconstrained reformulation and minimizing the objective $f(Q(x))$ with respect to $x\in\mathbb{R}^{d}$ (Li2017BinaryConnect). The key issue here is the non-differentiability of quantization operator $Q$. As such, the straight-through estimator of (bengio2013estimating) is the default choice to approximate the gradient of $f(Q(x)$. Specifically, $\nabla_{x}[f(Q(x))]=JQ(x)^{\top}\cdot\nabla f(Q(x))$ is approximated by $\nabla f(Q(x))$. Effectively, this replaces the Jacobian of the quantization operation by the identity matrix Li2017BinaryConnect; Hou2019QM; Chen2021QAdamEF.
@@
Error feedback reformulation. One way to gain more insight into the dynamics of STE is to examine it through the lens of error feedback. It is easy to observe that SGD-based training with STE over the iterate $x_{t}$ is equivalent to an instance of error-feedback chen2021quantizedadamerrorfeedback with quantized parameters $w_{t}=Q(x_{t})$ and quantization error $e_{t}=x_{t}-Q(x_{t})$. This equivalence is described over iterations $t\geq 0$ below:
$x_{t+1}=x_{t}-\alpha\widetilde{\nabla}f(Q(x_{t}))\iff\begin{cases}g_{t}=\alpha\widetilde{\nabla}f(w_{t})-e_{t}\par\\ w_{t+1}=Q(w_{t}-g_{t})\\ e_{t+1}=(w_{t}-g_{t})-w_{t+1}.\end{cases}$
@@
Effectively, the right-hand-side formulation says that STE is equivalent to a process where the quantization error $e_{t+1}$ due to weight quantization at a given step $(w_{t}-g_{t})-w_{t+1}$ is fed into the gradients at the next step. This feedback loop would be “exact”, fully correcting the error, if the loss function $f$ were an isotropic quadratic function, i.e. with identity Hessian, but is not the right update in general.
Multi-objective perspective. Prior work on convergence guarantees for the problem described in Equation (1) in the non-convex setting attempts to find a quantized point $Q(x^{*})$ that is a stationary point for the loss $f$, namely $\nabla f(Q(x^{*}))=0$.
However, as quantization is not an invertible transformation, this goal cannot be achieved in general, and current convergence guarantees have non-vanishing terms in their rate, proportional to quantization error lin2020dynamicmodelpruningfeedback; chen2021quantizedadamerrorfeedback; panferov2025unifiedscalinglawscompressed.
@@
For the sake of illustration, consider a toy example where our scalar loss function is $f(x)=\frac{1}{2}(x-\frac{1}{2})^{2}$, and the quantization operator $Q(x)=\lfloor x\rfloor$ gives the integer part of the one dimensional input $x\in\mathbb{R}$. Clearly, $\nabla f(Q(x))=\lfloor x\rfloor-\frac{1}{2}$ does not vanish at any point, with the smallest absolute value being $\frac{1}{2}$, coming from the quantization/rounding error.
@@
Instead, we propose to view the QAT problem in Equation (1) from the perspective of multi-objective optimization. The first objective is to minimize the loss $f(x)$: in the non-convex setting this becomes finding a stationary point, namely solving $\nabla f(x)=0$. The second objective is to satisfy the quantization constraint $x\in{\mathcal{Q}}$. In other words, we aim to find a solution $x^{*}$ for which $\nabla f(x^{*})=0$ and the distance between $x^{*}$ and $Q(x^{*})$ is minimized. As noted before, it is easy to see that these two conditions do not have to hold in general, meaning $\nabla f(x)$ can be non-zero for any quantized point $x\in{\mathcal{Q}}$. Instead, we define solutions as Pareto optimal points that balance these two conditions.
@@
Let $x\in\mathbb{R}^{d}$ be the current state of parameters. If we update towards $\nabla f(x)$ (or any other direction that aligns with it) with small enough step-size, then we would minimize the loss. Clearly, the opposite direction would increase the loss. Analogously, if we update towards Q(x)xQ(x)-x, then we would reduce quantization error. To have those two objectives in balance, we want our iterates to converge to some Pareto-optimal state $x^{*}$ such that $\nabla f(x^{*})=\lambda(Q(x^{*})-x^{*})$ for some scalar $\lambda>0$. In this case, any sufficiently small update applied to $x^{*}$ would hurt at least one of the objectives. In other words, any local improvement of one objective would be at the cost of other. Therefore, we define $\lambda$-Pareto optimal solution $x^{*}$ for the problem (1) as
@@
where $\lambda>0$ is a parameter balancing the two objectives. Recalling the toy example mentioned earlier, the value $x^{*}(\lambda)=\frac{1}{2(1+\lambda)}$ is $\lambda$-Pareto optimal for any $\lambda>0$. Note that there can be many Pareto optimal solutions for the same $\lambda$. In our toy example, $x^{*}=-\nicefrac{{1}}{{4}}$ is another $1$-Pareto optimal value.
Motivated by this multi-objective perspective and by the new optimality condition in Equation (2), we propose incorporating the quantization error $e_{t}=x_{t}-Q(x_{t})$ directly into the training dynamics. By doing so, we implicitly regularize the initial task loss, which can be made explicit if the quantization scheme is smooth (as we demonstrate in the theoretical analysis). Indeed, under the smoothness assumption of $Q$ (Assumption 3), the quantization error can be expressed as $x-Q(x)=\nabla\phi(x)$ for some smooth regularizer $\phi(x)$. Notably, our approach naturally avoids the need to handle the (non-existent) Jacobian $JQ(x)$ of the quantization operator and bypasses the computation of a proximal step for the explicit “quadratic” regularizer $\frac{\lambda}{2}\|x-Q(x)\|_{2}^{2}$, which would be infeasible for the dynamic quantization used in our experiments bai2018proxquant.
Error coupling.
@@
If the base optimizer is SGD, then these two corrections are identical. We will provide theoretical convergence guarantees to a Pareto-optimal point in the smooth non-convex setting. However, for stateful optimizers such as Adam, these two corrections produce conceptually different schemes, which we discuss below (subsuming bias correction terms into the learning rate $\alpha$):
Quantization error and mini-batch gradient
$\qquad e_{t}=x_{t}-Q(x_{t})$
$\qquad g_{t}=\widetilde{\nabla}f(x_{t}){\color[rgb]{0,0.88,0}\;+\;\lambda_{t}e_{t}},\;{\color[rgb]{0,0.88,0}\textrm{(coupled correction)}}$
Optimizer states
$\qquad m_{t}=\beta_{1}m_{t-1}+(1-\beta_{1})g_{t},$
$\qquad v_{t}=\beta_{2}v_{t-1}+(1-\beta_{2})g_{t}^{2},$
Optimizer update
$\qquad\Delta_{t}=m_{t}/(\sqrt{v_{t}}+\varepsilon){\color[rgb]{0,0,1}\;+\;\lambda_{t}e_{t}},\;{\color[rgb]{0,0,1}\textrm{(decoupled correction)}}$
Model update
$\qquad x_{t+1}=x_{t}-\alpha\Delta_{t},$
To compare these two correction variants in the context of the Adam optimizer, note that the coupled correction term $\color[rgb]{0,0.88,0}\lambda_{t}e_{t}$} becomes $(1-\beta_{1}){\color[rgb]{0,0.88,0}\lambda_{t}e_{t}}/(\sqrt{v_{t}}+\varepsilon)$ after being processed by the optimizer states. In contrast to the decoupled correction, the coupled version introduces diagonally-preconditioned curvature-aware correction term, where the curvature is approximated using the Adam statistics already maintained within the optimizer states.
@@
Concurrent work kwun2025lotion proposes an approach called LOTION, which smooths the quantized loss using unbiased randomized rounding (RR) via $f(Q(x))\approx\mathbb{E}_{\epsilon\sim{\rm RR}}[f(x+\epsilon)]$, and considers its second-order expansion as a surrogate loss to minimize, i.e., $f(x+\epsilon)\approx f(x)+\epsilon^{\top}\nabla f(x)+\tfrac{1}{2}\epsilon^{\top}{\bf H}(x)\epsilon$. Due to the unbiasedness of rounding (i.e., $\mathbb{E}[\epsilon]=0$), the first-order term of the expansion vanishes in expectation, and the smoothed loss becomes the original loss $f(x)$ with an additional curvature-aware regularization term $\tfrac{1}{2}{\rm Tr}\left({\bf H}(x)\,{\rm Cov}[\epsilon]\right)$. Here, the exact Hessian matrix is replaced by its Gauss–Newton approximation. If we further apply the empirical Fisher approximation and approximate the diagonal entries of the Gauss–Newton matrix using Adam statistics, this regularization takes similar form as our coupled correction variant (except the square root). Exactly computing the gradient estimator for the regularized loss would require differentiating the regularizer, which would involve third-order derivatives of the loss and is impractical. As such, this step is proposed to be skipped by the authors.
In contrast to LOTION, we propose to regularize the training dynamics directly, rather than the loss itself. One advantage of this approach is that the regularized dynamics naturally arise from our Pareto-optimality condition, while the practical benefit is that it avoids the need to differentiate or compute the proximal operator of a loss regularizer. In addition, as described in their paper, LOTION requires a full-precision forward pass, and the final model is a weight-only quantized model. By contrast, CAGE training is done with weights and activations quantized, where it can also benefit from high-performance low-precision GEMM kernels for more efficient training and inference.